MATH 135 – Calculus 1 Practice Problems on Evaluating Limits September 23, 2013

Directions

“Pair up” and work through the following questions. We’ll compare notes at the end of the period.

1. Does the limit lim_{x→1}(x^{2} + 5x + 3)^{2} exist? If so, find it and say how you are using the Limit Laws to do this.

2. Does the limit lim_{t→−3} \frac{t^{2}−9}{2t^{2}+7t+3} exist? If so, find it.

3. Same question for lim_{y→0} (\frac{1}{y} − \frac{1}{y^{2}+y} ).